If a swimming pool has a perimeter of 132 feet and is 21 feet wide, what is the area?
Answer by amarjeeth123(569)

Let the length be x.
Perimeter=2(length+width)
2(x+21)=132
x+21=66
x=45 feet
The length is 45 feet.
Area=45(21)=945 square feet
The area is 945 square feet.